# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ V SOTA in AI/ML-Assisted Positioning
As discussed above, AI/ML-assisted positioning leverages AI technologies for estimating positioning-related parameters, which are subsequently used to enhance the performance of traditional positioning algorithms. This approach is necessary because, while conventional estimation techniques are effective under ideal conditions, they often face significant challenges in real-world environments, including multipath propagation, signal distortion, and hardware imperfections. The advent of ML and DL offers new possibilities for addressing these challenges by employing data-driven methods to model the inherent complexity and nonlinearity of wireless environments. ML and DL techniques not only improve the robustness of parameter estimation but also enhance real-time performance in dynamic scenarios. Based on these advancements, we have conducted a detailed survey of AI-based Positioning Parameter Estimation algorithms, with a specific focus on AI-based LOS/NLOS detection, TOA/TDOA estimation, and angle estimation algorithms.
@@ V-A AI-based LOS/NLOS Detection
The wireless environment significantly impacts positioning accuracy. LOS scenarios, with unobstructed transmitter-receiver paths, yield more accurate distance and angle measurements, while NLOS scenarios, involving reflections and obstructions, degrade reliability due to multipath effects and delays. Thus, identifying LOS and NLOS conditions is critical for ensuring positioning accuracy.
@@
Prior to the advancements in deep learning, significant work had already been done in this area, employing ML algorithms such as support vector machines (SVM), Random Forest, Gradient Boosting Decision Tree [183], and AdaBoost for LOS/NLOS identification. The authors in [184] compare the performance of these traditional ML algorithms in LOS/NLOS identification tasks. To address the challenges in labeled data collection, [185] applied an unsupervised ML approach called expectation maximization for Gaussian mixture models to classify LOS and NLOS components.
Recent advancements in AI technology have significantly enhanced the capability to detect these conditions. In [186], the authors compared three ML classifiers, i.e. SVM, Random Forest, and MLP, to identify LOS and NLOS in UWB systems. In [187], in vehicle-to-vehicle (V2V) scenarios, the authors used the power angular spectrum (PAS) and employed SVM, Random Forest, and ANN for LOS/NLOS classification.
Subsequently, more advanced models and algorithms have been widely studied for LOS/NLOS identification. In [188], the authors propose a CNN-based neural network using PAS. In [189], a reversible transformation method is proposed for denoising CIR data, employing CNN to identify NLOS signals. The authors in [190] propose a Morlet wavelet transform and CNN approach for NLOS identification in indoor UWB positioning. In [191], the paper proposes a LOS/NLOS identification algorithm based on one-dimensional wavelet packet analysis and CNN. To fully utilize the characteristics of wireless channels, the authors in [192] use resource block (RB)-level eigenmatrix and eigenvector, designing a neural network based on 3D CNN for NLOS detection. To reduce computational complexity, the authors in [193] use both manually extracted features and features from a CNN based on raw CIR, designing a lightweight network with only 7.39% of the computational complexity compared to traditional CNNs. Similarly, in [194], the authors design a CNN-based lightweight LOS/NLOS recognition model with CIR as input and achieved good accuracy. In [195], the authors use CNN and GRU for NLOS identification based on CIR. In substation scenarios [196], the paper designs a neural network based on self-attention to separately learn the information contained in manually extracted channel features and the original channel features to enhance identification performance. The authors in [197] propose a dual-channel neural network with the consideration of the time and time-frequency domain characteristics. In [198], the paper proposes an algorithm based on Stockwell transform and CNN, utilizing transfer learning to enhance the algorithm’s performance. Additionally, the authors in [199] propose a learning method based on CNN, Bidirectional LSTM, and transfer learning for NLOS detection. In UWB scenarios [200], the authors propose a robust NLOS identification method based on transfer learning and GAN, utilizing cross-domain mapping from source and target domains to construct representative homogeneous features for both domains. In mmWave scenarios [201], the paper models the problem as a semi-supervised anomaly detection problem by exploiting a deep autoencoding kernel density model to identify several key parameters describing sparse spatiotemporal channel responses. Additionally, the authors in [202] perform LOS/NLOS identification based on RSS and RTT instead of CIR.
@@
Furthermore, various works have leveraged LOS/NLOS identification to improve positioning algorithms. For instance, the authors in [203] utilize LSTM to estimate approximate NLOS errors, correcting the positioning estimation results to enhance positioning accuracy. In UWB scenarios, the paper [204] converts CIR to Gramian angular fields (GAF) and uses CNN for NLOS identification, greatly enhancing ranging accuracy and reducing positioning errors based on identification results. In 5G channels, the authors in [205] use the angle-delay channel response matrixs for NLOS identification and select different positioning methods based on the identified LOS/NLOS results to enhance positioning accuracy.
@@ V-B AI-based TOA/TDOA Estimation
As mentioned above, TOA and Time TDOA are fundamental techniques in wireless positioning, providing the basis for accurate positioning by measuring the time it takes for a signal to propagate between transmitters and receivers. However, traditional TOA/TDOA estimation methods face significant challenges in real-world environments, including the presence of multipath propagation, NLOS conditions, and hardware imperfections. Recent advancements in AI can effectively enhance TOA/TDOA estimation by leveraging data-driven learning techniques.
@@
Early studies applied traditional ML algorithms to TOA estimation, focusing on simple models to mitigate errors in distance measurements [206, 207]. For example, the authors in [206] propose a SVM approach for error mitigation in UWB ranging systems. In [207] the authors use kernel principal component analysis for TOA estimation. In [208], the authors utilise SVM and semi-supervised learning to achieve ranging estimation. While these methods provided initial improvements, their inability to effectively interpret high-dimensional channel features limited their robustness and adaptability in complex environments.
To overcome the limitations of traditional ML, deep learning techniques have been employed to enhance TOA/TDOA estimation. These methods leverage the ability of neural networks to model complex, nonlinear relationships in data, achieving superior accuracy [209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223]. For example, in [209], the authors introduce an ANN-based RTT estimator for WiFi. The authors in [210] utilize ANN and radial basis function (RBF) networks to output distance errors rather than direct positions, enhancing TOA measurement accuracy. In addition, the authors in [211] propose error-mitigating ANNs to further refine TOA/TDOA estimates. In [212], the paper uses deep CNNs to exploit WiFi preamble structures for TOA estimation, effectively handling multipath environments. Furthermore, the paper proposes a novel dictionary filtering method leveraging neural networks for denoising and compressive sensing for channel impulse response extraction [213]. In [214], the authors use neural networks to generate high-resolution CIR for accurate TOA estimation. In the 5G system, the authors propose a CNN-based algorithm for TOA estimation [215]. In the IoT system, the authors in [224] propose a CNN-based algorithm by generating fine-grained features from full-band and resource-block-based reference signals, leveraging spectrogram-like cross-correlation feature maps with ML to directly project time-frequency domain variations into TOA results. In [216], the authors combine neural networks with Kalman Filters for TOA estimation using 5G downlink signals. In addition, the paper [217] addresses the challenge of false peaks in UL transmissions of LTE-LAA with a CNN-based TOA estimator tailored for the Block Interleaved Frequency Division Multiplexing (B-IFDM) structure. Considering hardware imperfections, the authors in [218] propose a CNN-based TOA positioning algorithm calibrated with CIR. Furthermore, considering the time dependence of CIR, the paper [221] proposes a BiRNN for accurate TOA estimation from CIR and an MLP for trilateration positioning. In addition, there are some works that consider the joint estimation of AOA and TOA to achieve higher accuracy estimation [225, 226, 227].
@@
In addition to the fundamental approach of end-to-end training with neural networks, recent research also leverages advanced AI techniques to address challenges such as data scarcity [113, 208, 228] and robustness [229] in TOA/TDOA estimation. To solve the difficulty of obtaining datasets, in [113], the paper combines neural networks with the Fine Timing Measurement (FTM) protocol to enhance WiFi RTT ranging. This work introduces an unsupervised learning framework that uses naturally accumulated sensor data to reduce data collection overhead. The authors in [228] utilize sensor data gathered during regular application use to reconstruct device trajectories. The reconstructed data is used to design an unsupervised learning technique for TOA estimation. The authors in [208] propose a semi-supervised learning approach for NLOS identification and mitigation, leveraging self-training with unlabeled measurements to achieve a 94% identification probability and reduce ranging error by 10%. To improve the robustness of the algorithm, the authors [229] propose inter-instance variational autoencoders (IIns-VAEs), which use variational inference with latent variables to simultaneously estimate distance and identify environmental conditions, demonstrating strong performance on real-world datasets.
@@ V-C AI-based Angle Estimation
In the era of traditional ML, algorithms such as Random Forest [230], SVM [231] and Regression Trees [232] have been widely employed for angle estimation. For example, the authors in [230] present an Random Forest-based optimization algorithm for AOA positioning, using weighted error mitigation to improve accuracy in NLOS environments. In [231], the paper presents an SVM-based AOA estimation method for vehicular communications. Although these methods are effective to some extent, they usually perform poorly in complex environments where multipath and NLOS effects dominate.
Compared to traditional ML techniques, neural networks have been employed due to their superior fitting capabilities, enabling more accurate angle estimation. Some studies demonstrate the potential of neural networks for angle estimation based on DNN [233, 234, 235, 236]. For instance, in [233], a deep learning-based framework integrates DNNs for efficient channel and DOA estimation in massive MIMO systems. To achieve super-resolution AOA estimation, the authors in [232] firstly obtain the MUSIC spectrum and then regard it as the input of the DNN, which increases the AOA estimation performance. In [236], the paper proposes a neural network-based method to reconstruct full-array covariance matrices from subarray samples, enabling improved DOA estimation with MUSIC. For device-free localization, the authors in [237] propose a learning-based AOA estimation, using a classifier to identify multipath components and a multilayer perceptron to correct AOA errors.
@@
With the development of neural network models, the accuracy of angle estimation has been further improved based on CNN [238, 239, 240, 80, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250], LSTM [251], Transformer [252, 253] and other advanced models[254]. The authors in [255] propose a deep convolutional network leveraging sparsity priors for efficient and precise DOA estimation, offering near real-time performance and improved robustness at low SNR. In [238], the paper introduces a deep ensemble learning approach for 2D DOA estimation, combining multiple independently trained CNNs to map spatial covariance matrices to azimuth and elevation angles to achieve angle estimation. In [239], the authors demonstrate the importance of phase features to improve DOA estimation accuracy, and propose a neural network framework, including DNN, 1-D CNN, and 2-D CNN with zero-padding for DOA estimation. Considering low-SNR conditions, the authors in [240] present a CNN-based method for robust DOA estimation, framing the problem as a multi-label classification task using sample covariance matrices. In [80], the authors introduce RFDOA-Net, a CNN-based DOA estimation method for UAV localization using nonuniform linear antenna arrays. In [241], the authors propose SDOA-Net, offering robustness to array imperfections and scalability to any target number with faster convergence. The authors in [242] introduce MoD-DNN, a model-driven DNN for AOA estimation, reformulating it as spatial spectrum image reconstruction, combining a CNN with a sparse conjugate gradient algorithm for automatic phase error calibration. Considering that the channels in wireless systems are complex, the authors in [243] introduce a complex-valued deep learning framework, and use virtual covariance matrices to handle spherical wave effects and complex signal features for near-field DOA estimation. Then, in [244], the paper employs a complex neural network-based deep learning approach and a parameterized algorithm for joint AOA and AOD estimation, enhancing computational efficiency via channel matrix preprocessing and coarse timing estimation. In addition, the authors in [251] propose an LSTM-based DOA estimation method that enhances phase features to improve accuracy and robustness to array imperfections. In [252], the authors propose a Transformer-based signal denoising network with temporal attention to enhance AOA estimation accuracy in indoor NLOS environments. The authors in [253] propose a Transformer-based sliding symbol detection method for ISAC, enabling simultaneous symbol detection, AOA and time delay estimation. Additionally, in [254], the paper proposes a cascaded neural network for DOA estimation of closely spaced sources, incorporating an SNR classification network and specialized estimation subnetworks for improved accuracy under low SNR and limited snapshots.
@@
Furthermore, advancements in AI technologies have further enhanced the performance of angle estimation algorithms. Autoencoders, known for their powerful feature extraction capabilities, have been increasingly applied in this domain. For example, in [256], the paper introduces a DNN framework combining autoencoders and parallel classifiers to achieve robust DOA estimation. This framework effectively addresses array imperfections and demonstrates strong generalization to unseen scenarios. In addition, transfer learning can improve the robustness of the DOA algorithm through domain transfer. The authors in [40] propose a transfer learning method using a ResNet for AOA estimation in massive MIMO systems, leveraging shared features across channel models to reduce data requirements and avoid training separate networks for each channel. Similarly, in [227], the authors propose a deep transfer learning-based DOA and TOA joint estimation algorithm using a multi-task network with shared-private structure, enabling efficient fine-tuning across different SNR scenarios for improved accuracy and reduced complexity. In the vehicle positioning system [257], the authors perform 2D DOA estimation for incoherently distributed sources with massive MIMO, employing dual 1D-CNNs, transfer learning, and attention mechanisms for improved accuracy, robustness, and efficiency.

### v2

@@ V SOTA in AI/ML-Assisted Positioning
TABLE VII: Summary of LOS/NLOS Classification Works
Backbone Algorithm Input Feature Key Insight
Traditional ML GMM [204], GBDT [205], k-NN, SVM, GMM, K-means [206] Signal energy, delay statistics [204], CIR statistics [205, 206] Handle simple features, rely on manual extraction from wireless channels.
FCNN Supervised learning [207, 208] CIR statistics [207, 208] Provide stronger feature extraction than traditional ML, but still rely on manually extracted features.
CNN Supervised learning [155, 209, 210, 211, 212, 213, 214, 215] CIR [155, 209, 211, 215], PAS images [210], Morlet-transformed CIR [212], Wavelet-packet images [213], CSI eigenmatrix/vector [214], and GAF [215] Handle high-dimension complex channel features, reduces manual feature engineering via deep feature extraction.
Transfer learning [140, 216] CIR [209], Stockwell Spectrograms [140] Enable cross-scenario generalization and adaptation.
GAN [216] CIR [216] Enhance model robustness by generating diverse training data.
Autoencoder [217, 218] ADCPM [217, 218] Learn latent representations of channel features to improve classification performance.
Temporal Model Supervised learning [219, 220, 221] CIR [219, 220], and CGWT [221] Capture temporal dependencies in CIR sequences using temporal models.
Transfer learning [222] CIR [222] Further enable cross-scenario generalization and adaptation.
Transformer Supervised learning [223, 224] CIR [224], CIR + statistics [223] Leverage self-attention mechanisms to fuse spatial and statistical features.
As discussed above, AI/ML-assisted positioning leverages AI technologies for estimating positioning-related parameters, which are subsequently used to enhance the performance of traditional positioning algorithms. This approach is necessary because, while conventional estimation techniques are effective under ideal conditions, they often face significant challenges in real-world environments, including multipath propagation, signal distortion, and hardware imperfections. The advent of ML and DL offers new possibilities for addressing these challenges by employing data-driven methods to model the inherent complexity and nonlinearity of wireless environments. ML and DL techniques not only improve the robustness of parameter estimation but also enhance real-time performance in dynamic scenarios. Based on these advancements, we conduct a detailed survey of AI-based positioning parameter estimation algorithms, with a specific focus on AI-based LOS/NLOS detection, TOA/TDOA estimation, and angle estimation algorithms.
@@ V-A AI-based LOS/NLOS Detection
The wireless environment significantly impacts positioning accuracy. LOS scenarios, with unobstructed transmitter-receiver paths, yield more accurate distance and angle measurements, while NLOS scenarios, involving reflections and obstructions, degrade reliability due to multipath effects and delays. Thus, identifying LOS and NLOS conditions is critical for ensuring positioning accuracy. We surveyed existing works in this area and summarized them in Table. VII.
@@
Prior to the advancements in deep learning, significant work had already been done in this area, employing ML algorithms such as support vector machines (SVM) [207], random forest [207], gradient boosting decision tree (GBDT) [205], Gaussian mixture models (GMM) [204], and k-means [206] for LOS/NLOS identification. With the development of AI technology, FCNN networks have also been widely used in this task. In [207] and [208], the authors compare three ML classifiers, i.e. SVM, random forest, and FCNN, to identify LOS and NLOS.
Subsequently, more advanced models and algorithms have been widely studied for LOS/NLOS identification. These works leverage deep learning backbones such as CNN [155, 209, 210, 211, 212, 213, 214, 215, 140, 216, 217, 218], temporal models (including LSTM [222, 220], GRU [219], and TCN [221]), and Transformer-based architectures [223, 224] to enhance classification performance. To enrich the input feature space and improve model discrimination capability, a variety of representations have been proposed, such as power angular spectrum (PAS) [210], wavelet-based transformations [212, 213], eigen features [214], Gramian angular fields (GAF) [215], Stockwell spectrograms [140], angle-delay cannel power matrix (ADCPM) [217, 218], and complex gaussian wavelet transform (CGWT) [221]. Furthermore, advanced learning techniques have been introduced to improve performance. In [140, 216] and [222], transfer learning is employed to enhance cross-scenario generalization and reduce the overhead of retraining models for new environments. In [216], GAN is employed to generate representative CIR samples for the target domain, effectively reducing the need for extensive data collection and manual labeling. Autoencoders [217, 218] provide a powerful means for unsupervised feature extraction and dimensionality reduction. They learn compact latent representations that effectively distinguish LOS from NLOS paths. In parallel, to reduce computational complexity and enable real-time deployment, several studies have proposed lightweight network architectures [155, 209, 214] without sacrificing much accuracy.
@@
Furthermore, various works have leveraged LOS/NLOS identification to improve positioning algorithms. For instance, the authors in [220] utilize LSTM to estimate approximate NLOS errors, correcting the positioning estimation results to enhance positioning accuracy. In UWB scenarios, the paper [215] uses CNN for NLOS identification, greatly enhancing ranging accuracy and reducing positioning errors based on identification results. In 5G channels, the authors in [218] use ADCPM for NLOS identification and select different positioning methods based on the identified LOS/NLOS results to enhance positioning accuracy.
@@ V-B AI-based TOA/TDOA Estimation
As mentioned above, TOA and TDOA are fundamental techniques in wireless positioning, providing the basis for accurate positioning by measuring the time it takes for a signal to propagate between transmitters and receivers. However, traditional TOA/TDOA estimation methods face significant challenges in real-world environments, including the presence of multipath propagation, NLOS conditions, and hardware imperfections. Recent advancements in AI can effectively enhance TOA/TDOA estimation by leveraging data-driven learning techniques.
TABLE VIII: Summary of TOA/TDOA Estimation Works
Backbone Algorithm Input Feature Key Insight
Traditional ML SVM [225, 226, 227], PCA [228], KNN [227] Channel statistics (signal amplitude, energy, delay spread, etc.) [225, 226, 228, 227] Handle simple features, relies on manual extraction from wireless channels.
FCNN Supervised learning [229, 230] CSI [229] and CIR [230] Provide stronger feature extraction than traditional ML, but still relies on manually extracted features.
CNN Supervised learning [231, 232, 233, 234, 154, 235, 236, 237, 238] CIR [233, 235, 236, 237, 238], channel cross-correlation [231, 234, 154] and IQ sample [232] Handle high-dimension complex channel features, reduces manual feature engineering via deep feature extraction.
VAE [239] CIR [239] Enable cross-scenario generalization and adaptation.
Temporal Model Supervised learning [240] CIR [240] Capture temporal dependencies in CIR sequences using temporal models.
Transformer Supervised learning [241, 167] CIR [241, 167] Leverage self-attention mechanisms to fuse spatial and statistical features.
@@
Early studies applied traditional ML algorithms to TOA estimation, focusing on mitigating errors in distance measurements [225, 228, 226]. For example, the authors in [225, 226] propose an SVM approach for error mitigation for ranging estimation. In [228], the authors use kernel principal component analysis (PCA) for TOA estimation. In [227], the authors combine ML algorithms with Kalman Filters for TOA estimation using 5G downlink signals. While these methods provide initial improvements, their inability to effectively interpret high-dimensional channel features limits their robustness and adaptability in complex environments.
To overcome the limitations of traditional ML, deep learning techniques have been employed to enhance TOA/TDOA estimation. These methods leverage the ability of neural networks to model complex, nonlinear relationships in data, achieving superior accuracy [229, 230, 231, 233, 232, 234, 235, 236, 237, 240, 238, 239, 241, 167]. These works leverage neural network backbones such as FCNNs [229, 230], CNNs [231, 232, 233, 234, 154, 235, 236, 237, 238, 239], temporal models [240], and Transformers [241, 167]. The related works are summarized in Table. VIII. For traditional ML-based methods, input features are typically derived from channel statistics such as signal amplitude, energy, and delay spread. In contrast, deep learning approaches can directly learn from raw or minimally processed channel information, including CIR, CSI, cross-correlation signals, and even raw in-phase and quadrature (IQ) samples. With the development of model architectures, neural networks are increasingly capable of modeling and extracting high-dimensional representations from these raw inputs through end-to-end training. This enables more robust feature learning and improves the generalization of TOA/TDOA estimation models in complex wireless environments. For example, in [233], the authors use neural networks to generate high-resolution CIR for accurate TOA estimation. In the 5G system, the authors propose a CNN-based algorithm for TOA estimation [232]. In the IoT system, the authors in [154] propose a CNN-based algorithm by generating fine-grained features from full-band and resource-block-based reference signals, leveraging spectrogram-like cross-correlation feature maps to directly project time-frequency domain variations into TOA results. In addition, some studies [242, 243, 244] consider joint angle and TOA estimation to achieve higher accuracy to achieve higher accuracy estimation.
@@
In addition to the fundamental approach of end-to-end training with neural networks, recent research also leverages advanced AI techniques [239] to address challenges such as data scarcity and robustness in TOA/TDOA estimation. In [239], the authors propose inter-instance variational autoencoders (VAEs), which use variational inference with latent variables to simultaneously estimate distance and identify environmental conditions. Furthermore, to solve the difficulty of obtaining datasets, in [77], the paper combines neural networks with the Fine Timing Measurement (FTM) protocol to enhance RTT ranging. This work introduces an unsupervised learning framework that uses naturally accumulated sensor data to reduce data collection overhead. The authors in [245] utilize sensor data gathered during regular application use to reconstruct device trajectories for designing an unsupervised learning technique for TOA estimation.
@@ V-C AI-based Angle Estimation
TABLE IX: Summary of Angle Estimation Works
Backbone Algorithm Input Feature Key Insight
Traditional ML SVM [246] PDP and path loss [246] High-dimensional channel of the antenna array makes it difficult to deploy traditional ML algorithms.
FCNN Supervised learning [247, 248, 249, 250, 251] Received signal [247] and covariance matrix [248, 249, 250], and spectrum vector from MUSIC [251]. Leverage the nonlinear fitting capabilities of CNNs to improve performance.
Autoencoder [252] covariance matrix [252] Learn latent representations of channel features to improve performance.
CNN Supervised learning [253, 254, 255, 51, 256, 257, 258, 259, 260, 261, 153, 262, 263] CIR [260, 261], CSI [259], covariance matrix [253, 255, 254, 257, 256, 258], correlation matrix [153, 262], and received signal [51, 263] Handle high-dimension complex channel features for angle estimation.
Transfer learning [264, 265, 244] covariance matrix [264], correlation matrix [265], and received signal [244] Enable cross-scenario generalization and adaptation.
Temporal Model Supervised learning [266] Covariance matrix [266] Capture temporal dependencies of the channel using temporal models.
Transformer Supervised learning [267, 268] Received signal [267, 268] Leverage self-attention mechanisms to capture long-term dependencies of channels.
The related works for angle estimation are summarized in Table. IX. In the era of traditional ML, algorithms such as SVM [246] have been employed for angle estimation. For example, in [246], the paper presents an SVM-based AOA estimation method for vehicular communications using power PDP and path loss received by each element of the antenna array. However, angle estimation inherently relies on antenna arrays, resulting in high-dimensional input features that are often difficult to process effectively using traditional ML algorithms. This complexity limits the performance and scalability of traditional methods in practical scenarios.
@@
Compared to traditional ML techniques, neural networks have been employed due to their superior fitting capabilities, enabling more accurate angle estimation. Some studies demonstrate the potential of neural networks for angle estimation based on FCNN [247, 248, 249, 250]. For instance, in [247], a deep learning-based framework integrates FCNNs for efficient channel and DOA estimation in massive MIMO systems. To achieve super-resolution AOA estimation, the authors in [251] firstly obtain the MUSIC spectrum and then regard it as the input of the FCNN, which increases the AOA estimation performance. In [250], the paper proposes a neural network-based method to reconstruct full-array covariance matrices from subarray samples, enabling improved DOA estimation with MUSIC. With the development of neural network models, the accuracy of angle estimation has been further improved based on CNN [253, 254, 255, 51, 256, 257, 258, 259, 260, 261, 153, 262, 263], LSTM [266], Transformer [267, 268]. For example, in [253], the paper introduces a deep ensemble learning approach for 2D DOA estimation, combining multiple independently trained CNNs to map spatial covariance matrices to azimuth and elevation angles to achieve angle estimation. In [254], the authors demonstrate the importance of phase features to improve DOA estimation accuracy, and propose a neural network framework for DOA estimation. Considering low-SNR conditions, the authors in [255] present a CNN-based method for robust DOA estimation, framing the problem as a multi-label classification task using sample covariance matrices. Considering that the channels in wireless systems are complex, the authors in [258] introduce a complex-valued deep learning framework, and use virtual covariance matrices to handle spherical wave effects and complex signal features for near-field DOA estimation. In [259], the paper employs a complex neural network-based deep learning approach and a parameterized algorithm for joint AOA and AOD estimation, enhancing computational efficiency via channel matrix preprocessing and coarse timing estimation. In addition, the authors in [266] propose an LSTM-based DOA estimation method that enhances phase features to improve accuracy and robustness to array imperfections. Additionally, in [267], the authors propose a Transformer-based signal denoising network with temporal attention to enhance AOA estimation accuracy in indoor NLOS environments.
@@
Furthermore, advancements in AI technologies have further enhanced the performance of angle estimation algorithms. Autoencoders, known for their powerful feature extraction capabilities, have been increasingly applied in this domain. For example, in [252], the paper introduces a DNN framework combining autoencoders and parallel classifiers to achieve robust DOA estimation. This framework effectively addresses array imperfections and demonstrates strong generalization to unseen scenarios. In addition, transfer learning can improve the robustness of the DOA algorithm through domain transfer. The authors in [265] propose a transfer learning method using a ResNet for AOA estimation in massive MIMO systems, leveraging shared features across channel models to reduce data requirements and avoid training separate networks for each channel. Similarly, in [244], the authors propose a deep transfer learning-based DOA and TOA joint estimation algorithm using a multi-task network with shared-private structure. In the vehicle positioning system [264], the authors perform 2D DOA estimation for incoherently distributed sources with massive MIMO, employing transfer learning and attention mechanisms for improved accuracy, robustness, and efficiency.
V-D Lessons Learned
The transition from traditional ML to deep learning has significantly improved positioning accuracy, particularly in complex wireless environments characterized by multipath propagation, NLOS conditions, and dynamic interference. Deep learning demonstrates strong capabilities in capturing high-dimensional channel features and modeling nonlinear relationships, thereby reducing the need for manual feature engineering. To further improve performance, the choice of input features is very important. While early approaches rely heavily on manually extracted channel statistics, more recent studies have shown the benefits of using raw or minimally processed inputs such as CIR, CSI, and received signal matrices, which preserve richer spatial and temporal information for learning. However, robustness across domains remains a significant challenge. Although many deep models perform well in specific scenarios, their performance often degrades in unseen environments. To address this issue, techniques such as transfer learning, domain adaptation, and data augmentation have emerged as promising solutions for improving generalization and reducing reliance on large labeled datasets.